A market consists of two individuals. Their demand equation are 

Q1 = 16 - 4P and Q2 = 20 – 2P, respectively. 


= a) What is the market demand equation?


b) At a price of $2, what is the point price elasticity for each person


1
Expert's answer
2021-02-04T14:46:42-0500

a) Market demand = Sum of individual demand

"MD = Q_1 + Q_2 \\

= 16 -4P + 20 -2P \\

= 36 -6P"

b) At P = $2

Point price elasticity:

"PPE = \frac{dMD}{dp} \times \frac{p}{MD} \\

PPE_{market} = \frac{d(36-6p)}{dp} \times \frac{2}{36-6p} \\

= -6 \times \frac{2}{36 – 6 \times 2} \\

= \frac{-12}{24} \\

= -0.5 \\

PPE_{individ.1} = \frac{d(16-4p)}{dp} \times \frac{2}{16-4p} \\

= -4 \times \frac{2}{16-4 \times 2} \\

= \frac{-8}{8} \\

= -1 \\

PPE_{individ.2} = \frac{d(20-2p)}{dp} \times \frac{2}{20-2p} \\

= -2 \times \frac{2}{20-2\times 2} \\

= \frac{-4}{16} \\

= \frac{-1}{4} \\

= -0.25"